OpenStudy (anonymous):

4^3 means to multiply 3 factors of 4?

Parth (parthkohli):

4^3 means to multiply 4 by itself 3 times.

OpenStudy (anonymous):

how 4^2 means 4*4, the same principal is applied only with three instead, 4*4*4

Parth (parthkohli):

\[\large4 ^3 = 4\times 4 \times4= 64 \]

OpenStudy (anonymous):

It simply means MULTIPLYING 4, 3 TIMES..

OpenStudy (anonymous):

So the statement is wrong?

OpenStudy (mathstudent55):

The statement is true. You multiply together three factor of 4, so you only perform two multiplications. 4^3 = 4 x 4 x 4 = 16 x 4 = 64

OpenStudy (anonymous):

You think yourself about this, we have given the hints..

OpenStudy (anonymous):

What to you is a 'factor'?

OpenStudy (anonymous):

For instance.....3 x 4 = 12. The factors are 3 and 4. The product is 12.

OpenStudy (anonymous):

Factors are the numbers you multiply to get the product